OpenStudy (anonymous):

(MEDALS!!!!) Please answer. Find the probabilities. Write your answer as a fraction, as a decimal, and as a percent. If necessary, round decimals to the nearest thousandth and percents to the nearest tenth. A game requires 12 red markers, 12 white markers, 12 purple markers, 12 yellow markers, and 12 green markers. What is the probability of randomly drawing a yellow marker?

OpenStudy (anonymous):

12/60 = 0.2 = 20%

OpenStudy (anonymous):

@sunnysunshine 12/60 as a fraction? I thought it was 1/5 or 2/5..

OpenStudy (anonymous):

it can simplify 12/60 = 1/5

OpenStudy (anonymous):

Oh, forgot to simplify. I always do that...